NEWS

These are the news starting 01 Jan 2006
Click here for the news of the previous years.

28 Dec 06 Main computer breakdown at 15:30. Repaired the 29th Dec. No data lost.
22 Dec 06 Maintenance and recalibration work... the last one for this year!
14 Dec 06 Main logger computer broke down on 12Dec06 12:00; no data lost due to buffering in logger. Correct operation resumed today at 11:00 UTC; graph will be ok from tomorrow on...
09 Dec 06 Recheck of the CO2 sensor: zero still ok, calibration factor slightly adjusted.
08 Dec 06 Maintenance and recalibration by Envitec. The CO2 sensor has been fully revised: cleaning of filters and chamber, recalibration etc.
03 Dec 06 Uploaded October and November 2006 data into archive. Sorry for the delay!
29 Nov 06 New recalibration of CO2 sensor.
27 Nov 06 I guess the problem with the CO2 sensors comes from a valve that became stuck during the cslibration process in the zero-gas position, so that calibration gas was mixed with zero gas and as such gave a reading too low. Around 00:00 that valve sprang back, so that the real gas readings were now too high (see steep rise). Calibration was repeated today using only the sample inlet; it seems that we are back to normal operation. As a matter of fact, the remarks of 25/11 on disregarding the Nov. CO2 data does probably not apply, as the malfunction began only during the calibration process on 25/11.
26 Nov
06
1. WBGT sensor has been switched off (actually the water of the wet thermometer has not been refilled) since a month ago, as heatstress-measurements are somewhat superfluous during the colder season. To make this more clear, a notice "Not measured" has been put on the WBGT icons. Nevertheless the dry and globe temperatures are still measured and are present in the data file.
2. There surely is a problem with the CO2 sensor. Please stay tuned....
25 Nov
06
Maintenance and calibration work starting 15:00. Ther was (is) a problem with the Mir9000 CO2 sensor: the measurement of span gas was much too low (420 instead of 496ppm); this has been corrected now by adjusting the calibration factor. A check with an Edinburg Sensor instruments now shows good according values. I'm not sure when this down-drift started since the last calibration at 22 Oct 06. The data from November should be taken with caution. More on this later....
22 Oct
06
There has been a computer problem from 19th to 20th Oct.. This has only an impact on the graphs and ASCII files at the live data screen. As everything has been stored on the logger, no data have been lost, and the monthly oct06.dat file (when issued) will be complete.
21 Oct 06 1. Uploaded Sep.06 data into data archive
2. Maintenance and recalibration on CO2 and Ozone sensors ( --> spikes!)
20 Oct
06
New short note on meteoLCD CO2 measurements by F. Massen
1 Oct 06 Main power loss between 16:30 and 17:00
05 Oct 06 Uploaded to the data archive: 051006.mt5 file with all total ozone column measurements by Microtops #5375 sun photometer from 11 June 2005 to 05 October 2006 (original, non edited Microtops file) .
03 Oct 06 The ground ozone sensor had a curious tendency to read negative O3 concentrations at night: an offset of 10 ug/m3 has been programmed today at 06:16 to take care of this.
28 Sep 06 Maintenance work and recalibration by Envitec.
25 Sep 06 Repaired ultrasonic sensor back in action starting 13:00; graphs switched from old cup anemometer back to ultrasonic at 19:30
23 Sep 06 Uploaded Aug.06 data into data archive.
16 Sep 06 Maintenance work and recalibrations.
12 Sep 06 Computer problem solved; live data available again (there is a small glitch on the previous days graphs...it is self-healing!)
11 Sep 06 Serious computer problem starting 09 Sep 22:30; no live data for the moment!
29 Jul 06 Total ozone column will be updated more sparingly, due to operator unavailability. During my leave, Frank Thillen will make regular measurements if possible using the older Microtops. These data will be adjusted to the new instrument readings using about 2 months of dual measurements. The complete data will be put online for mid-September.
11 Jul 06 New paper on UVR and ozone trends by F. Massen et al.
05 Jul 06 Uploaded Jun 06 data
02 Jul 06 Restarted a new all.dat file (thaht's the file you get by the link "All files since last logger restart" on the page with the live data. The file all_01Jul05to30Jun06. has been put into the data archive
30 Jun 06 Maintenance work by Envitec.
26 Jun 06 My face is RED! We had a very impressive thunderstorm yesterday from about 13.00 to 18 UTC; I tried to make a 1-second-rate AVI file; a power loss due to lightning corrupted this file, and, worse, the regular 1-minute pictures that are the basis of the mpg movies, have not been made during the avi-recording (reason unknown). Sorry, you wont see much of that storm on the 20060626.mpg movie....
24 Jun 06 Pre-release of a rather important paper by F. Massen & N. Harpes on ozone and UVB trends measured from 1998 to 2005 at meteoLCD
13 Jun 06 The ultrasonic anemometer (air_speed, ch.11) went crazy starting 04 Jun 06 06:00. Windspeed data now come from the older mechanical cup anemometer (ch.61). If the problem is solved, we will switch back. 
Thx to Nico Harpes for his sharp eyes!
02 Jun 06 Uploaded April and May 06 data (apr06.dat, may06.dat) into data archive
26 May 06 Maintenance and recalibration work. Discard big CO2 and NO swings during today.
20 May 06 break down of communications with web server due to an UPS failure at about 03:30. UPS replaced and operation resumed at 08:30. No data loss.
17 May 06 Marcel Kramer has written a very nice meteoLCD widget for Mac OSX users. Bravo Marcel!
13 May
06
Installed air conditioning unit in room where the gas sensors are located; the Logger_Temp column in the Ascii data file shows a drop of about 4C in room temp.
12 May 06 Changed link to SST (Sea Surface Temperature) anomalies to a (better) web site displaying all measured SST's and allowing to pick past data for viewing.
11 May 06 New booklet on good practice for sun exposure from the Ligue Medico Sociale of Luxembourg. German and French version.
08 May 06 1. A research team from the Indian Institute of Tropical Meteorology has published a research paper on precipitable water based on meteoLCD Microtops data. Look here for an abstract!
2. Fan of CO2 sensor replaced (spike in data is due to power down)
02 May 06 New intercomparison report of total ozone data between Uccle, Diekirch and 2 satellites
28 Apr 06 Maintenance by Envitec
26 Apr
06
meteoLCD has joined www.blitzortung.org, a free non commercial net of 40 lightning tracking stations of Western and Central Europe; realtime lightning data are compiled from the participant stations and are available free of charge.
25 Apr 06 Reactivated WetBulbGlobeTemperature (WBGT) heatstress sensor for this season: new wick, some cleaning, and 2 times a day automatic refill with distilled water
22 Apr 06 Upload of March06 data files.
06 Apr 06 New pump installed. Ground ozone data from 03Apr 07:00 to 06Apr 16:00 possibly 10 ug/m3 low.
New proxy computer for webcam installed.
04 Apr 06 The main pump for the air manifold has broken down. A temporary new sampling line was installed for the O3 sensor; back to correct readings at 15:00. A new pump has been ordered.
03 Apr 06 Ozone sensor malfunction starting 07:00. Will be repaired asap.
26 Mar 06 1. There was a problem with the lightning moviemaker script. Corrected (thx to Gilles Massen)
2. Changed several plots to display horizontal gridlines (if one single or two same axis)
17 Mar 06 Calibration and maintenance work.
12 Mar 06 I added two neoworkx counters showing the countries on-line and where the past visitors came from. This might be interesting; if you find these counters annoying, please tell me so, or just give me your opinion. f.m.
07 Mar 06 Gamma radiation sensor again online. Defunct computer replaced.
05 Mar 06 Uploaded feb06.dat data file into data archive.
27 Feb 06 1. I think the problem with the intermittent system failure has been found (thanks to Gilles Massen from Restena). A ftp process did not release its resources, so open processes slowly accumulated up to a point where the Linux server making the plots was choked.
2. A couple of weeks the computer reading the gamma radiation sensor broke down. I did not have the time to repair this, but it will be done asap.
17 Feb 06 There was a software problem with the lightning page, so that the radar screen was not updated. Anyway, no storms happened, and everything is working correctly since 20:00 today.
13 Feb 06 Main Linux computer replaced, and much needed housekeeping done. As always, when you fiddle with working software, the surprises come at the next startup. I think most glitches have been solved by now.
10 Feb
06
Full data for 2005 uploaded into archive. The yearly edited and commented files are being repacked into individual files, like 2005_only.xls, as the big Excel file has become unusable (it is still there for the years 1998 to 2004). See more in the data archive.
08 Feb
06
Computer problem Graphs somewhat disturbed!
17 Jan 06 Maintenance and calibration work.
07 Jan
06
Uploaded dec05.dat and 2005.zip files into data archive. The 2005.zip file is the concatenation of all 12 monthly data files (including headers with comments) of the year 2005.
03 Jan 06 The usual new year problem with the 7 days graphs.. as the logger file does not contain the year, Gnuplot gets confused the 1rst week of every new year. This will be corrected the 7th January.

 


Back Home